### v1

# Textual Similarity as a Key Metric in Machine Translation Quality Estimation

Abstract
Machine Translation (MT) Quality Estimation (QE) assesses translation reliability without reference texts. This study introduces “textual similarity” as a new metric for QE, using sentence transformers and cosine similarity to measure semantic closeness. Analyzing data from the MLQE-PE dataset, we found that textual similarity exhibits stronger correlations with human scores than traditional metrics (hter, model evaluation etc.). Employing GAMMs as a statistical tool, we demonstrated that textual similarity consistently outperforms other metrics across multiple language pairs in predicting human scores. We also found that “hter” actually failed to predict human scores in QE. Our findings highlight the effectiveness of textual similarity as a robust QE metric, recommending its integration with other metrics into QE frameworks and MT system training for improved accuracy and usability.
Keywords: translation quality evaluation, sentence transformers, cosine similarity, regression analysis, key metric
1 Introduction
There are two kinds of quality evaluations in machine translation (MT): “MT quality evaluation” and “MT quality estimation” [11]. Each field has distinct standards and applications, ensuring comprehensive translation quality assessment. The former is used to compare the quality of translated texts when the reference translations are available. In contrast, the latter, also named as MT “quality estimation (QE)”, involves assessing the quality of MT outputs without reference translations [8]. QE is essential for various applications, such as determining whether an automatically translated sentence or document is ready for the end user or requires human post-editing. It can flag passages with critical errors, serve as a quality metric when reference translations are unavailable, and assist in computer-aided translation interfaces by highlighting text needing human revision and estimating the required human effort. The current study focuses on the latter, quality estimation.
“MT quality evaluation” involves using various automatic metrics to assess translation accuracy and effort. Common metrics include BLEU, which measures n-gram overlap with reference translations and applies a brevity penalty; METEOR, which considers precision, recall, synonyms, and stemming; TER, which calculates the number of edits needed to match a reference translation; chrF, which evaluates character n-grams, useful for morphologically rich languages; and BERTScore, which uses pre-trained embeddings to assess semantic similarity. Moreover, supervised metrics, trained using human-judged data often show a higher correlation with human evaluations, and some metrics were used such as BEER, BLEND [6].
However, in MT QE, several features play pivotal roles in both quality estimation and enhancing the overall translation process. The “model eveluation” (ML_eval) quantifies the MT model’s confidence in its translations. This metric is typically derived from the log-likelihood of the translation given the source sentence. Higher scores reflect greater confidence, indicating that the translation aligns well with the model’s learned patterns from the training data. These scores are crucial for quality estimation TM systems, helping to predict the reliability of translations from TM models and highlight areas that may require further human intervention or correction [2], [4].
Another significant feature is “human translation edit rate” (hter) [10], and it measures the number of edits required to correct a translation output from MT to match a human reference. It accounts for insertions, deletions, substitutions, and shifts, providing a quantitative measure of the effort needed for post-editing. This makes hter a practical metric for assessing the quality of MT systems, as it directly correlates with the human effort required to produce accurate translations.
In MT quality estimation, “ML_eval” often serve as one key feature in helping to gauge the reliability and quality of TM systems when the reference translations are not available. The metric of “ML_eval”, combined with linguistic features and contextual embeddings, enhance the accuracy of TM quality assessments. They help identify translations that may require human review or post-editing. Further, in the training or fine-tuning of NMT (neural machine translation) or LLMs-based TM models, the metric of “ML_eval” still plays an important role, such as on optimizing translation performance by leveraging large parallel corpora, word alignments, and contextual learning techniques [5], [7]. Even if the metric is not directly used in training or fine-tuning, it still could provide valuable feedback on the translation quality, guiding iterative improvements and fine-tuning of the model. This distinction underscores the dual role of “ML_eval” in both evaluating and enhancing translation quality.
While “ML_eval” provides valuable insights into QE, the metric has certain limitations. One significant issue is the tendency to overestimate confidence in poor translations, often due to overfitting or a lack of diverse training data. This overconfidence can lead to inaccuracies in QE. Moreover, the metric is sensitive to the training data distribution, and meaning biases or gaps in the data can skew the scores. Third, this metric may also fail to capture contextual meaning and subtle semantic differences that human evaluators can detect, leading to discrepancies between model scores and perceived translation quality.
“hter” has also some weaknesses. First, the variability in human editing styles and preferences can lead to inconsistent hter scores, making it challenging to standardize quality assessments. Additionally, “hter” focuses on the number of edits rather than the nature of the changes, failing to differentiate between minor stylistic tweaks and substantial corrections. This can skew the perception of translation quality. Furthermore, the metric relies on the availability of high-quality reference translations, which may not always be accessible, limiting its applicability in certain contexts. These limitations suggest that while “hter” is useful, it should be complemented with other evaluation methods for a more comprehensive assessment of translation quality.
We have identified some limitations in the two key metrics used in Quality Estimation (QE). Despite these, traditional QE methods face a significant, often overlooked problem: the reliance on correlation statistical analysis for evaluation. Correlation is primarily used to understand the relationship between two data sets. Correlation only indicates the degree to which two sets of data are related, but it does not establish a relationship between them. However, to determine the effect of one variable on another, regression analysis is essential. For a deeper insight into key features affecting machine learning quality estimation data, regression analysis provides a more robust approach. For example, if we want to determine whether “hter” has an effect on human scores, regression analysis is necessary. The correlation between “hter” and human scores does not provide this level of insight.
To address these weaknesses, it is crucial to complement existing metrics with additional features, continuously update and diversify training data, and incorporate new features to enhance the reliability and usability of QE. Upgrading statistical analysis methods is also necessary to gain deeper insights. This study proposes that “textual similarity” could be considered a key feature in MT quality estimation and training. Using existing MT quality estimation datasets, we employed advanced statistical methods to compare this new metric with existing features and explore their advantages.
2 Methods
2.1 Dataset
MLQE-PE is a comprehensive dataset for MT QE and Automatic Post-Editing, covering eleven language pairs, including both high- and low-resource languages [4]. The dataset features up to 10,000 translations per language pair, annotated with sentence-level direct assessments, post-editing effort, and word-level binary good/bad labels. Each source-translation pair includes the post-edited sentence, article titles, and details of the neural MT models used. The dataset is thoroughly documented and analyzed, and the information on the baseline system performances is included.
The main variables in the MLQE-PE includes: original (original sentence), translation (MT output), scores (list of DA [direct assessment] scores by all annotators - the number of annotators may vary), mean (average of DA scores), z_scores (list of z-standardized DA scores), z_mean (average of z-standardized DA scores), model_scores (NMT model score for sentence). The other information on sentence translation such as “hter” (human translation edit rate) is also included in the dataset. Note that “model_scores” is the the MT model’s confidence in its translation. According to MLQE-PE, these MT systems are SOTA ones, representing the recent advancements in MT models. In this way, “model_scores” signifies a direct indication of the translation’s reliability from MT, that is, the translation quality assessment. The higher scores represents the higher translation quality. Conversely, a lower “hter” indicates that a translation needs more human efforts to post-edit, that is, the translation quality is higher. Moreover, we also created some factors. For instance, the standard variation (sd) of different human scores for the same translation, and the human annotator number.
There are 11 language pairs available: English-German (en-de), English-Chinese (en-zh), Romanian-English (ro-en), Estonian-English (et-en), Nepalese-English (ne-en), Sinhala-English (si-en), and Russian-English (ru-en).  [^1] Although the paper on MLQE-PE claimed that their dataset also includes other language pairs, the dataset on these language pairs in the github they provided is not available.
2.2 Textual similarity
Textual similarity (or sentence similarity) measures how similar or different two pieces of text are semantically. This can involve comparing sentences, paragraphs, or entire documents to see how closely they match in meaning or content. It can be calculated using various techniques, such as comparing word overlaps, semantic similarity, or using machine learning models to assess how alike the texts are. HHowever, after transformers revolutionized deep learning, textual similarity can now be effectively measured using existing language models. [12], [3], [13]
Pretrained sentence transformers are highly effective for generating sentence or textual embeddings due to their ability to capture deep semantic meanings [9]. By leveraging the Transformer architecture, these models encode sentences into high-dimensional vectors that reflect their contextual and semantic content, rather than merely their lexical features. This allows for accurate comparisons of semantic similarity between sentences or texts, which is crucial for tasks like paraphrase detection, information retrieval, and clustering.
The “sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2” model creates dense vector representations of sentences, capturing the semantic essence of the text. These vectors are placed in a shared embedding space where similar meanings are mapped to nearby points. The model’s fine-tuning for tasks like paraphrase identification ensures that sentences with similar meanings are close in this space. The use of sentence transformers is particularly valuable in multilingual contexts, as they support multiple languages and capture semantic nuances effectively. The task in the current study is to understand the source text in one language and the translated texts in the other language. The capability of multilingual sentence transformer models is essential for applications to explore the semantic similarity among languages. The current study employs this multilingual sentence transformer model to compute semantic similarity for one source sentence and its translation. Our fundamental approach involves using the sentence transformer to generate text embeddings and then calculating their similarity using the cosine method.
2.3 Regression statistical analysis
We applied Generalized Additive Mixed Models (GAMM) in analyzing how factors influence human scores. GAMMs incorporate non-linear relationships between the dependent and independent variables through smooth functions [14]. GAMMs allow for both fixed and random effects, accommodating complex variations within hierarchical data structures. The “additive” part of GAMM means that the model expresses the dependent variable as a sum of smooth functions of predictors, along with any random effects and an error term. This flexibility makes GAMMs particularly useful for modeling non-linear trends in data, where the effect of variables is not strictly linear and may vary by group or over time.
GAMM could leverage the function s(). This smooth function better gets model fittings for some factors, and the interaction smooth could find the interaction among some given factors. Some random variables could play a very important role, such as different language pairs in TM. The role of such random variables could be well explored by using GAMMs. GAMMs are friendly to make model comparison by referring to AIC (Akaike information criterion is an estimator of prediction error).
In GAMM setups, the independent variable is “human score”, and other metrics are dependent variables. Some random factors, such as human evaluator number, different language pairs (source language - target language), could play an essential role [1]. A GAMM fitting should include a number of factors. The reason for this is that various metrics or factor could predict human score or take effect on human score, and these factors co-work to play a role. The purpose of using GAMM fitting is to better explore how these factors take effects on human scores.
3 Results
3.1 Result 1: Correlations
We plotted the Pearson correlations among various factors, as shown in Fig. LABEL:fig. The results indicate that “ML_eval” is correlated with the human score (mean) at 0.15, with “hter” at 0.06, and with the human score (z-mean) at 0.3. “Textual similarity” shows a correlation of 0.47 with the human score (mean), -0.06 with the human score (z-mean), -0.45 with “hter”, and -0.21 with “ML_eval”. Overall, “textual similarity” exhibits stronger correlations with other factors compared to “ML_eval”.
Figure 1: correlations among various factors
3.2 Result 2: Overall performance
We established three types of GAMM fittings, where “human score (mean)” is the independent variable, and other factors include “ML_eval”, “textual similarity”, “sd of human scores”. “human evaluator number”, “language pairs (langs)” are incorporated as random variables. We chose “human score (mean)” over “human score (z-mean)” as the independent variable primarily because “human score (mean)” closely follows a normal distribution. The GAMM setups are detailed below, and the results are presented in Table 1.
  • base model = bam(human score $\sim$ s(ML_eval)+s(textual similarity)+s(sd)+
    s(hter)+ s(evaluator num, bs=“re”)+s(langs, bs=“re”))
  • m1=bam(human score $\sim$ s(ML_eval)+s(sd)+s(hter)+s(evaluator num, bs=“re”)
    +s(langs, bs=“re”))
  • m2=bam(human score $\sim$ s(textual similarity)+s(sd)+s(hter)+s(evaluato num, bs=“re”)+s(langs, bs=“re”))
  • m3=bam(human score $\sim$ s(ML_eval)+s(textual similarity)+s(sd)+s(evaluator num, bs=“re”)+s(langs, bs=“re”))
Here, the base model demonstrates the best performance when all factors are included. However, each of the remaining models includes only a subset of these factors. The Akaike Information Criterion (AIC) is used to represent the performance of this GAMM fitting. The base model has the lowest AIC. When the AIC of model $m1$ is subtracted from the AIC of the base model, the resulting $\Delta$AIC indicates the contribution of “textual similarity”, as “m1” does not include “textual similarity” compared to the base model. A smaller $\Delta$AIC also indicates better performance and greater contribution for a given factor. The results are shown in Table 1
Table 1: $\Delta$AIC for different GAMM fittings with random smooths. A smaller $\Delta$AIC indicates better performance (n=45886)
factor (contribution) $\Delta$AIC
(m1- base model) contribution of textual similarity 848.87
(m2- base model) contribution of ML_eval 8785.82
(m3-base model) contribution of hter 586.3
The model with “textual similarity” outperforms the one with “ML_eval”. However, the model with “hter” shows the best performance overall. This is an overall trend. Despite this, it is important to examine the performance across individual language pairs. The next section explores the performance of these factors in each language pair.
3.3 Result 3: Across language pairs
Using the similar GAMM setups in the previous section, we explore how these factors perform across seven language pairs. For each language pair, we established the same GAMM fittings (i.e., the base model, the model without textual similarity, the model without “ML_eval”, and the model without “hter” ). $\Delta$AIC is the subtraction of a specific model from the base model. Note that the random variable “ language pairs” is not included because we implmented GAMM fittings for an indiviual language pair among the seven pairs. After carrying out four GAMM fittings for each of seven language pairs, we plotted the partial effects for each factor of interest, as shown in Figure 2. When the p-value is greater than 0.05, the plot is not significant. Within the same language pair, a smaller $\Delta$AIC indicates better performance. Clearly, all cases for “ML_eval” are significant, and there is only one insignificant case for “textual similarity”. In contrast, “hter” has only two significant cases, indicating five insignificant cases.
Compared with “ML_eval”, “textual similarity” outperforms in the cases of “German-English”, "English-Chinese”, “Romanian-English”, “Russian-English” and “Sinhala-English”. Generally, “textual similarity” demonstrates the best performance across individual language pairs.
Figure 2: The partial effects on human score from different factors. The x-axis represents the specific metric being analyzed, while the y-axis indicates human score. Each curve within a plot illustrates the relationship between a predictor variable (plotted on the x-axis) and the response variable. Steeper slopes on these curves indicate a stronger influence of the predictor variable on human score. Conversely, gentler slopes imply a weaker influence, indicating that changes in the predictor variable have a less pronounced effect on human score. Such plots could give deep insights on the relationship between one given metric and human score. The information on p-value and $\Delta$AIC is shown at the top of each plot.
4 Discussion
Considering the correlation, we found that “textual similarity” exhibits a strong correlation with human scores, “ML_eval”, and “hter”. In contrast, “ML_eval” shows a weaker correlation with the human score (mean). When examining the GAMM fittings overall and across various language pairs, the performance of “textual similarity” surpasses that of “ML_eval”. On an individual language pair basis, “textual similarity” also outperforms both “ML_eval” and “hter”. In short, “textual similarity” as a metric could be closely correlated with human score, and is highly capable of predicting human score.
The following provides a detailed analysis of each metric. “ML_eval” demonstrates a significant impact across all language pairs, indicating that this metric is both useful and effective for QE. However, “hter” does not show a significant impact on human scores in the most cases, suggesting that this metric may not be suitable for evaluating MT quality in certain language pairs. On the other hand, “textual similarity” predicts human scores in the majority of language pairs, highlighting its potential as an effective metric for QE.
The present study underscores the importance of selecting appropriate metrics for QE. “Textual similarity”, in particular, emerges as a robust and reliable metric that consistently correlates with human evaluation scores and effectively predict human scores, making it a valuable tool for improving the accuracy of QE. In contrast, while “ML_eval” remains a useful metric, its effectiveness varies across different language pairs. “hter”, despite being commonly used, may not be sufficient in some cases, necessitating the consideration of alternative or supplementary metrics like textual similarity.
It is easy to understand why textual similarity is such an effective metric. When a translated text is semantically close to the source text, it indicates that the translation meets a crucial standard of quality: the meaning of the translation should closely match the original text. In contrast, the “hter” metric often fails for most language pairs because post-editing efforts do not necessarily reflect changes in meaning. For example, a translation may be very close in meaning to the original text but contain some misspellings. Humans may still consider the translation to be good, even though the post-edit rate is low due to the necessary corrections. Conversely, minor edits to core verbs or key words in a translation text may require minimal edit effort but significantly alter the meaning. However, such a translation text does not necessarily meet the standard of good translation. This discrepancy shows that post-edit rate does not always correlate with translation quality.
Therefore, it is worthwhile to explore the cognitive recognition of what constitutes a good translation for different language users. Understanding this can significantly enhance both machine translation and quality estimation processes by aligning them more closely with human judgments of translation quality.
To date, “textual similarity” has not been proposed as a metric in QE. Our study applied a variety of statistical methods to analyze data from established QE datasets, demonstrating that this metric is both reliable and effective compared to commonly used metrics in QE. Our findings provide strong evidence that “textual similarity” is a robust metric, making it a valuable addition to the existing suite of QE metrics. Given its proven effectiveness, “textual similarity” should be included as a key metric in QE and incorporated into the training of machine translation (MT) systems. This integration can enhance the accuracy and reliability of MT quality assessments, ultimately improving the performance and usability of MT systems.